OpenStudy (anonymous):

Can someone please show me step by step how to solve the equation and check for extraneous solutions? 2|7-7x|=2x+4

OpenStudy (anonymous):

2(7−7x)=2x+4 I will try to solve your equation. −14x+14=2x+4(Simplify both sides of the equation) −14x+14+−2x=2x+4+−2x(Add -2x to both sides) −16x+14=4 −16x+14+−14=4+−14(Add -14 to both sides) −16x=−10 (−16x) (−16) = (−10) (−16) (Divide both sides by -16) x=0.625

OpenStudy (anonymous):

Thank you... I understand it now ^-^

OpenStudy (anonymous):

No problem! :D

OpenStudy (anonymous):

You also should solve for -2(7-7x)=2x+4 since the absolute values is involved.

OpenStudy (anonymous):

You can check that \( x =\frac 3 2\) is a solution for your problem in addition to the root that @Nikkeyiboo found which was \( x =\frac 5 8\)

OpenStudy (anonymous):

So the answer as a whole would be

OpenStudy (anonymous):

x = 3/2 or x = 5/8 ?

OpenStudy (anonymous):

It may be. Convert the decimal into a percent, or if easier a fraction and you can get your answer.